OpenStudy (anonymous):

What percent of 23.6 is 59 ?

OpenStudy (lgbasallote):

\[R = \frac{P}{B}\] where R means rate P means percentage B is base here R is missing P is 59 B is 23.6 \[R = \frac PB\] \[R = \frac{59}{23.6}\] solve for the quotient then multiply by 100 to get the percent

OpenStudy (lgbasallote):

 it can help you if you know how to use it

OpenStudy (anonymous):

I got 2.5 So, is the final answer is 2.5% or 25% ?

OpenStudy (lgbasallote):

2.5 is right for 59/23.6 now multiply it by 100  that's done by moving two decimal places to the right